Exercise 11-10 Disposal of property, plant, and equipment [L011-2] Mercury Inc. purchased equipment in 2016 at a cost of $497,000. The equipment was expected to produce 580,000 units over the next five years and have a residual value of $33,000. The equipment was sold for $253,600 part way through 2018. Actual production in each year was: 2016 -83,000 units; 2017 133,000 units; 2018 67,000 units. Mercury uses units-of-production depreciation, and all depreciation has been recorded through the disposal date. Required 1. Prepare the journal entry to record the sale. 2. Assuming that the equipment was sold for $280,000, prepare the journal entry to record the sale.
Answer #1

Answer

  • Workings

A

Original cost

$           497,000.00

B

Salvage value

$              33,000.00

C = A - B

Depreciable base

$           464,000.00

D

Usage in units

                     580,000

E = C/D

Depreciation per unit

$                        0.80

F = 83000 + 133000 + 67000

No. of units produced till sale

                     283,000

G = E x F

Accumulated Depreciation till date of sale

$           226,400.00

  • Requirement 1

Accounts title

Debit

Credit

Cash

$           253,600.00

Accumulated Depreciation - Equipment

$           226,400.00

Loss on sale of Equipment

$              17,000.00

   Equipment

$                        497,000.00

  • Requirement 2

Accounts title

Debit

Credit

Cash

$           280,000.00

Accumulated Depreciation - Equipment

$           226,400.00

   Gain on sale of Equipment

$                             9,400.00

   Equipment

$                        497,000.00